OpenStudy (anonymous):

find the speed on the point on the rim of a 24 cm diameter fly wheel which is turning at 2800 revolutions per minute. give your answer in m/s

OpenStudy (anonymous):

given diameter = 24 cm so raius radius = 24/2 = 12cm =12/100 =0.12 m angular speed is given = 2800rev /min =2800* 2pi /60 rad /sec can you do it from here

OpenStudy (anonymous):

linear speed = angular speed * radius of rotation (u may use pi=22/7)

OpenStudy (anonymous):

just multiply ( 2800* 2pi /60 ) *12 m/s